Koontz Company manufactures a number of products. The standards relating to one of these products are shown below, along with actual cost data for May.

Standard Cost per Unit Actual Cost per Unit
Direct materials:
Standard: 1.80 feet at $1.60 per foot $

2.88

Actual: 1.75 feet at $2.00 per foot $ 3.50
Direct labor:
Standard: 0.90 hours at $18.00 per hour

16.20

Actual: 0.95 hours at $17.40 per hour 16.53
Variable overhead:
Standard: 0.90 hours at $6.00 per hour 5.40
Actual: 0.95 hours at $5.60 per hour 5.32
Total cost per unit $

24.48

$ 25.35
Excess of actual cost over standard cost per unit $ 0.87

The production superintendent was pleased when he saw this report and commented: “This $0.87 excess cost is well within the 5 percent limit management has set for acceptable variances. It's obvious that there's not much to worry about with this product."

Actual production for the month was 11,500 units. Variable overhead cost is assigned to products on the basis of direct labor-hours. There were no beginning or ending inventories of materials.

Required:

1. Compute the following variances for May:

a. Materials price and quantity variances.

b. Labor rate and efficiency variances.

c. Variable overhead rate and efficiency variances.

2. How much of the $0.87 excess unit cost is traceable to each of the variances computed in (1) above.

3. How much of the $0.87 excess unit cost is traceable to apparent inefficient use of labor time?
Answer #1

Part 1 a

material price variance = Actual quantity * (actual price –standard price) = 1.75*(2.00-1.60) = $0.70 U

Total material price variance = 11500*0.70 = $8050 U

Material quantity variance = Standard price *( actual quantity- standard quantity) = 1.60*(1.75-1.80) = $0.08 F

Total Material quantity variance = 11500*0.08 = $920 F

Part 1 b

Labor rate variance = Actual direct labor hours * (Actual labor hour rate - Standard labor hour rate) = 0.95*(17.40-18.00) = $0.57 F

Total Labor rate variance = 11500*0.57 = $6555 F

Labor efficiency variance = Standard price per hour * (Actual hours - Standard hours required for actual production) = 18.00*(0.95-0.90) = 0.9 U

Total Labor efficiency variance = 11500*0.90 = $10350 U

Part 1 c

Variable overhead rate variance = Actual direct labor hours * (Actual labor hour rate - Standard labor hour rate) = 0.95*(5.60-6.00) = 0.38 F

Total variable overhead rate variance = 11500*0.38 = $4370 F

Variable overhead efficiency variance = Standard price per hour * (Actual hours - Standard hours required for actual production) = 6.00* (0.95-0.90) = 0.3 U

Total variable overhead efficiency variance = 11500*0.30 = $3450 U

Part 2

Material price variance

$0.70 U

Material quantity variance

$0.08 F

Labor rate variance

$0.57 F

Labor efficiency variance

$0.90 U

Variable overhead rate variance

$0.38 F

Variable overhead efficiency variance

$0.30 U

Excess of actual over standard cost per unit

$0.87 U

Part 3

Labor efficiency variance

$0.90 U

Variable overhead efficiency variance

$0.30 U

Excess of actual over standard cost per unit

0.87 U

Less portion attributable to labor inefficiency:

0.90 U

Labor efficiency variance

0.30 U

Variable overhead efficiency variance

1.20 U

Portion due to other variances

$0.33 F

In case if there would not have been for the apparent inefficient use of labor time, the total variance in unit cost for the month would have been favorable by $0.33 rather than unfavorable by $0.87.